OpenStudy (richyw):

find a scalar potential energy function \(U(x,y,z)\) so that \[\nabla U = \left\langle \frac{-kx}{\left( x^2+y^2+z^2 \right)^{3/2}}\, ,\, \frac{-ky}{\left( x^2+y^2+z^2 \right)^{3/2}}\, ,\, \frac{-kx}{\left( x^2+y^2+z^2 \right)^{3/2}} \right\rangle\]

OpenStudy (richyw):

so I'm not sure how I could go about this without a huge amount of algebra. First of all I was wondering if I am allowed to factor the denominator out to show that the vector field is conservative like this...\[\nabla U = \frac{-k}{\left( x^2+y^2+z^2 \right)^{3/2}}\left\langle x\,,\,y\,,\,z\right\rangle\]and then\[\frac{\partial U_1}{\partial y}=0=\frac{\partial U_2}{\partial x}\]\[\frac{\partial U_1}{\partial z}=0=\frac{\partial U_3}{\partial x}\]\[\frac{\partial U_2}{\partial z}=0=\frac{\partial U_3}{\partial y}\] so \(\text{curl}\,\nabla U=0\)

OpenStudy (richyw):

After that is there any tricks I can use to save pages of writing when trying to figure this out?

OpenStudy (sirm3d):

\[\large \nabla U=<U_x, U_y, U_z>\]\[\large U_x=\frac{ -kx }{ (x^2+y^2+z^2)^{3/2} } \rightarrow U=\int\limits_{}^{} \frac{ -kx }{ (x^2+y^2+z^2)^{3/2} }dx +g(y,z)\] problem: \[\large \left| U \right|=\sqrt{U _x^2+U_y^2+U_z^2}\]

OpenStudy (sirm3d):

is the third component of \[\large \nabla U\]\[\large \frac{k\color{red}x}{(x^2+y^2+z^2)^{3/2}}\] really kx and not kz?

OpenStudy (richyw):

oops yes it is kz. typo

OpenStudy (richyw):

so I integrate that wrt x and then I know that my constant of integration is a function of y and z only right?

OpenStudy (richyw):

so now I take the partial of that function wrt y?

OpenStudy (richyw):

sorry i'm just doing that now but I have \[U=\frac{k}{\left( x^2+y^2+z^2\right)^{1/2}}+h(y,z)\] and I am going to find \(U_y\)

OpenStudy (sirm3d):

yep

OpenStudy (richyw):

which I get \[U_y=\frac{-ky}{\left( x^2+y^2+z^2\right)^{3/2}}+\frac{\partial g(y,x)}{\partial y}\]

OpenStudy (sirm3d):

or \[\large \frac{ \partial g }{ \partial y }=0\] which means g(y,z) is a function of z only

OpenStudy (richyw):

right ok got that.

OpenStudy (sirm3d):

computing Uz, we conclude that \[\large \frac{dg}{dz}=0\] so g(z) is a numeric constant

OpenStudy (richyw):

alright I just arrived there as well!

OpenStudy (sirm3d):

therefore \[\large U=\frac{ k }{ (x^2+y^2+z^2)^{1/2} }+C\] the scalar potential function. i was wrong earlier in \[\large \left| U \right|=\sqrt{U_x^2+U_y^2+U_z^2}\]

OpenStudy (richyw):

ok thanks a lot. I really appreaciate it. I understand how to do this clearly now!

OpenStudy (sirm3d):

YW